# The Tropical Geometry of Subtraction Schemes

### v1

@@ Abstract
We study the construction of local subtraction schemes through the lenses of tropical geometry. We focus on individual Feynman integrals in parametric presentation, and think of them as particular instances of Euler integrals. We provide a necessary and sufficient condition for a combination of Euler integrands to be locally finite, i.e. to be expandable as a Taylor series in the exponent variables directly under sign of integration. We use this to construct a local subtraction scheme that is applicable to a class of Euler integrals that satisfy a certain geometric property. We apply this to compute the Laurent expansion in the dimensional regulator $\epsilon$ of various Feynman integrals involving both UV and IR singularities, as well as to generalizations of Feynman integrals that arise in effective field theories and in phase-space integrations.
@@ 1 Introduction
where the “counterterm” $I^{\rm ct}(\epsilon)$ is a simpler (i.e. already known) integral with the same pole part as the desired one $I(\epsilon)$, so that the “renormalized” expression $I^{\rm ren}(\epsilon)$ is finite as $\epsilon\to 0$. We will be interested in promoting this understanding of UV and IR divergences to the integrand level. The motivation is two-fold. From a computational point of view, one may be interested in the direct evaluation of the finite quantity $I^{\rm ren}(\epsilon)$ by integrating, numerically or analytically, an appropriate renormalized integrand; rather than computing separately the quantities $I(\epsilon)$ and $I^{\rm ct}(\epsilon)$ and explicitly (i.e. analytically) exposing the cancellation of the poles in $\epsilon$. For this purpose, a result such as (1) is not enough, as we will illustrate with simple examples below. On a more conceptual level, one may hope that understanding subtraction schemes at the level of the integrand may unearth geometrical structures underpinning the physics of renormalization, in analogy to how positive geometries have given a new perspective on the origin of locality and unitarity of scattering amplitudes - and simplified their computation Arkani-Hamed:2013jha ; Arkani-Hamed:2017tmz ; Arkani_Hamed_2018 ; arkanihamed2023loop .
@@
Here we propose a different strategy, based on the development of a suitable subtraction scheme. We will explain how to build a collection of counterterms that reproduce the pole structure of an Euler integral, so that subtracting them from the original integrand yields a renormalized integrand that can be safely expanded under the sign of integration. Furthermore, the counterterm integrands can be integrated exactly, explicitly manifesting the pole structure of the Euler integral under consideration. In other words we look for a counterterm integrand $\mathcal{I}^{\rm ct}$ such that,
@@
$I(\epsilon)^{\rm ren}=\int_{0}^{1}\frac{dx}{x}\left[x^{\epsilon}(1+x)^{\epsilon}-x^{\epsilon}\right]=\int_{0}^{1}\frac{dx}{x}(\log(1+x)\epsilon+\dots)=\zeta(2)/2+\dots,$ (11)
@@
Our first major result is a necessary and sufficient condition for a combination of Euler integrands to be locally finite, Th. 3. Building on this, we develop a simple-minded subtraction scheme, inspired by ideas discussed in Brown:2019wna in the context of string amplitudes, and already considered in the context of UV divergent Feynman integrals Hillman:2023ezp . These earlier works focus on infinite families of physically relevant Euler integrals: tree-level string amplitudes and UV divergent Feynman integrals, respectively. On the other hand, our subtraction scheme is applicable to a class of Euler integrals that satisfy a certain geometrical property. Although in a sense this goes beyond the integrals considered in the aforementioned literature, it is also more restrictive: in particular, the subtraction scheme does not apply to the whole families of integrals treated in Brown:2019wna ; Hillman:2023ezp .
@@ 2 Polyhedral Geometry
A cone $\sigma$ in $\mathbb{R}^{n}$ is any set closed under taking positive spans. That is, if vectors $\{\rho_{i}\}_{i\in I}$ are in $\sigma$, then so is $\rm{Span}_{+}(\{\rho_{i}\}_{i\in I})$. A cone can be written as the positive span of a collection of vectors called rays,
$\sigma=\mathrm{Span}_{+}\ \mathrm{Rays\ }\sigma,$ (17)
@@
$\mathrm{Poly}\left(\left(\begin{array}[]{c}0\\ 0\\ 4\\ 4\\ 0\\ \end{array}\right),\left(\begin{array}[]{cccc}0&2&0&-1\\ 0&0&2&-1\\ 0&-2&0&-1\\ 0&0&-2&-1\\ 0&0&0&1\\ \end{array}\right);0,0\right),$
@@
$\mathcal{P}=\{0-\rho_{+}\cdot{\bf z}\leq 0\}\cap\{-\rho_{-}\cdot{\bf z}\leq 0\}\cap\{1-\rho_{e}\cdot{\bf z}=0\},$ (27)
@@ 3 Newton Polytopes
Together, the value of $\mathrm{Trop\ }P(\rho)$ and the initial form $P|_{\rho}$ control the leading behavior of $P$ under the rescaling $\alpha\to\alpha\lambda^{-\rho}$, in the limit $\lambda\to 0$, which motivates us to understand them better. It turns out that they are respectively piecewise linear and piecewise constant on the fan $\Sigma=\rm{Fan\ Newt\ }P$. To see this, consider a cone $\sigma\in\Sigma$. For any $\rho\in\sigma^{+}$ the functional $\rho\cdot{\bf z}$ is always maximized over $\mathrm{Newt\ }P$ at the same face $F_{\sigma}$ of the Newton polytope. This shows that $P|_{\rho}$ is constant as $\rho$ varies in $\sigma^{+}$. Furthermore, over $\sigma$ we can compute $\mathrm{Trop\ }P(\rho)$ as $\rho\cdot{\bf m}$ for any ${\bf m}$ on the face $F_{\sigma}$,
@@
Consider the polynomial $P=1+x+y+xy$. Its Newton polytope and normal fan are shown in Fig.2 (a) and (b). We have $P|_{(-1,0)}|_{(0,-1)}=(1+y)|_{(0,-1)}=1=P|_{(0,-1)}|_{(-1,0)}$. On the other hand, $P|_{(-1,0)}|_{(1,0)}=1+x\neq y(1+x)=P|_{(1,0)}|_{(-1,0)}$.
@@ 4 Euler Integrals
At this point it should be clear that understanding the facet presentation of the Newton polytope of the Symanzik polynomials $\mathcal{U}$ and $\mathcal{F}$, that appear in the parametric form of Feynman integrals, is a mathematical problem of great physical interest. Currently, this presentation is not available if not in special cases Schultka:2018nrs  [^5] See also Arkani-Hamed:2022cqe for conjectures beyond this case. In particular, this means that there is no known purely graph theoretical algorithm - i.e. one that avoids solving a linear programming problem - to establish the convergence of an arbitrary Feynman integral.
@@ 4.2 Sector Decomposition
Due to the positivity assumption on the coefficients of the polynomials, the only problem may arise by integrating around a neighborhood of the origin, where the integrand can be replaced by its series expansion. For each integrand $\mathcal{I}_{i}$, separate the Taylor series in (60) into the part $S_{i}$ with ${\bf m}$ on a hypecube with sides given by the integer part of the exponents, $\left\lfloor{\nu}\right\rfloor$, and the rest $\mathcal{R}_{i}=\mathcal{O}(t^{1+\left\lfloor{\nu}\right\rfloor})$. In order for $\mathcal{I}=\sum_{i}\mathcal{I}_{i}$ to be renormalized, it has to be that $\sum_{i}t^{\nu_{i}}S_{i}=0$. The derivative $\left.\frac{\partial^{k}}{(\partial\epsilon)^{k}}\mathcal{I}\right|_{\epsilon=0}$ is the coefficient of $\epsilon^{k}$ in the series
@@ 5 Tropicalization
$\rm{Trop\ }f(t)=\lim_{\alpha^{\prime}\to 0}\alpha^{\prime}\log f(exp(-t/\alpha^{\prime})).$ (65)
If we identify $t$ with the logarithm of $\alpha$, we can think of the tropicalization of a function $f$ as capturing the asymptotic behavior of $f$ at the boundaries of the hyperoctant $\mathbb{R}_{+}^{n}$.
@@
where $\tau_{\lambda}$ is the operator returning the standard Taylor series expansion in $\lambda$. In practice, to compute $\tau_{\rho}$ one factors from the rescaled integrand $\mathcal{I}(\alpha\lambda^{-\rho})$ the overall power $\lambda^{-\mathrm{Trop\ }\mathcal{I}(\rho)}$, then performs a Taylor expansion on the rest and finally multiplies back in the factor $\lambda^{-\mathrm{Trop\ }\mathcal{I}(\rho)}$ (in complete analogy with the operators $\tau_{z}$ described in the Introduction). The structure of the Puiseaux series is illuminated by the notion of initial form of the integrand:
@@
Let us now consider again an Euler integrand $\mathcal{I}$ and two vectors $\rho_{1}$ and $\rho_{2}$ that are compatible according to the structure of the fan $\Sigma$, i.e. that belong to a common cone. Then we have
@@ 5.1 Local Finiteness for Euler Integrals
Note that while the proofs of the above theorems are based on the idea of decomposing the integrals into sectors, of which there is essentially one per vertex of $\mathrm{Newt\ }\mathcal{I}$, the result is that one can check a condition on each facet of $\mathrm{Newt\ }\mathcal{I}$. In general, the number of vertices and of facets of a polytope can be vastly different. For instance, in the case of Feynman integrals [^8] With some caveat regarding special kinematics already mentioned in 5 the relevant polytopes are generalized permutohedra Schultka:2018nrs , which have roughly $E!$ many vertices against $2^{E}$ many facets, $E$ being the number of edges of the graph. Therefore, it is practically useful to know about the above theorems, rather than performing sector decomposition and then apply Th. 1 to the resulting sector integrals.
@@ 6 Subtraction scheme
We anticipate that the geometric property is in general not satisfied by generic Euler integrals, nor by the specific instances that show up in physics applications. Therefore, this assumption represents the fundamental limitation of our subtraction scheme. Even so, the geometric property is satisfied often enough for the result to be practically useful. We will substantiate this claim by virtue of examples in Section 7.
@@
The combinatorics behind the above formula is familiar from the work of BPHZ on UV renormalization Hepp1966 ; Zimmermann1969 , and similar forest formulae have been written down also for IR divergences CHETYRKIN1982340 ; CHETYRKIN1984419 ; Herzog:2017bjx . The novelty compared with these previous results is that the renormalization map defined here produces an integrand that is not only finite but also locally finite. This is the first main result of this section,
@@ 7.4 Double box
which agrees with the numerical values obtained by FIESTA.
@@ 7.5 Eye Graph
$\Gamma(3\epsilon)^{-1}I(s;\epsilon)=\frac{1}{\epsilon^{2}}+\frac{7}{\epsilon}+(31-6\zeta(2))+\mathcal{O}(\epsilon).$ (113)
@@ 7.7 An angular integral
We provide another example beyond the case of Feynman integrals, a class of phase-space integrals known as angular integrals . In smirnov2024expansion was presented a parametric formulation which allows to apply our master formula to this case as well. In the language of smirnov2024expansion , we consider the four denominators, one-mass integral, given by
@@
The integrand has a power-divergence along the ray $\rho=(0,0,0,1)$, which requires analytic continuation in that direction to reduce log-divergent integrands. This produces two integrands $\mathcal{I}_{1,2}$ satisfying the geometric property and therefore we can apply our subtraction formula. Adopting to the notation of Panzer_2015 for multiple polylogarithms $\mathrm{Li}_{\vec{n}}$, we find
$I_{1}=$ $-\frac{1}{8\epsilon(r_{+}-r_{-})}\Bigg{(}\log\left(\frac{r_{+}}{r_{-}}\right)\log\left(\frac{v_{14}v_{44}}{2v_{12}v_{24}}\right)-\text{Mpl}[\{2\},\{1+r_{+}\}]+\text{Mpl}[\{2\},\{1+r_{-}\}]$
$\quad-\text{Mpl}[\{1,1\},\left\{\frac{(1+r_{+})v_{13}}{v_{13}-v_{14}},\frac{v_{13}-v_{14}}{v_{13}}\right\}]+\text{Mpl}[\{1,1\},\left\{\frac{(1+r_{-})v_{13}}{v_{13}-v_{14}},\frac{v_{13}-v_{14}}{v_{13}}\right\}]$
$\quad-\text{Mpl}[\{1,1\},\left\{\frac{(1+r_{+})v_{23}}{v_{23}-v_{24}},\frac{v_{23}-v_{24}}{v_{23}}\right\}]+\text{Mpl}[\{1,1\},\left\{\frac{(1+r_{-})v_{23}}{v_{23}-v_{24}},\frac{v_{23}-v_{24}}{v_{23}}\right\}]$
$\quad+\text{Mpl}[\{1,1\},\left\{\frac{2(1+r_{+})v_{34}}{2v_{34}-v_{44}},\frac{2v_{34}-v_{44}}{2v_{34}}\right\}]-\text{Mpl}[\{1,1\},\left\{\frac{2(1+r_{-})v_{34}}{2v_{34}-v_{44}},\frac{2v_{34}-v_{44}}{2v_{34}}\right\}]\Bigg{)}+\mathcal{O}(\epsilon^{0}),$
with
$r_{\pm}=-\frac{v_{14}v_{23}+v_{13}v_{24}-v_{12}v_{34}\pm\sqrt{(v_{14}v_{23}+v_{13}v_{24}-2v_{12}v_{34})^{2}+2v_{13}v_{23}(v_{12}v_{44}-2v_{14}v_{24})}}{2v_{13}v_{23}},$
and
$I_{2}=$ $\frac{1}{\epsilon}\left(-\frac{\log\left(\frac{v_{14}}{v_{12}}\right)}{4v_{12}v_{13}v_{14}}+\frac{\log\left(\frac{v_{14}}{v_{13}}\right)}{4v_{12}v_{13}(v_{13}-v_{14})}-\frac{\log\left(\frac{v_{14}}{v_{13}}\right)}{2v_{12}(v_{13}-v_{14})v_{14}}-\frac{\log\left(\frac{v_{24}}{v_{12}}\right)}{4v_{12}v_{23}v_{24}}+\frac{\log\left(\frac{v_{24}}{v_{23}}\right)}{4v_{12}v_{23}(v_{23}-v_{24})}\right.$
$\left.-\frac{\log\left(\frac{v_{24}}{v_{23}}\right)}{2v_{12}(v_{23}-v_{24})v_{24}}-\frac{\log\left(\frac{v_{34}}{v_{13}}\right)}{4v_{13}v_{23}v_{34}}-\frac{\log\left(\frac{v_{34}}{v_{23}}\right)}{4v_{13}v_{23}v_{34}}\right)+\mathcal{O}(\epsilon^{0}),$
and finally
$I({\bf v};\epsilon)=8\pi\epsilon(I_{1}+I_{2})+\mathcal{O}(\epsilon).$ (122)

### v2

@@ Abstract
We study the construction of local subtraction schemes through the lenses of tropical geometry. We focus on individual Feynman integrals in parametric presentation, and think of them as particular instances of Euler integrals. We provide a necessary and sufficient condition for a combination of Euler integrands to be locally finite, i.e. to be expandable as a Taylor series in the exponent variables directly under sign of integration. We use this to construct a local subtraction scheme that is applicable to a class of Euler integrals that satisfy a certain geometric property. We apply this to compute the Laurent expansion in the dimensional regulator $\epsilon$ of various Feynman integrals involving both UV and IR singularities, as well as to generalizations of Feynman integrals that arise in effective field theories and in phase-space integrations, for which we provide new analytic results.
@@ 1 Introduction
where the “counterterm” $I^{\rm ct}(\epsilon)$ is a simpler (i.e. already known) integral with the same pole part as the desired one $I(\epsilon)$, so that the “renormalized” expression $I^{\rm ren}(\epsilon)$ is finite as $\epsilon\to 0$. We will be interested in promoting this understanding of UV and IR divergences to the integrand level. The motivation is two-fold. From a computational point of view, one may be interested in the direct evaluation of the finite quantity $I^{\rm ren}(\epsilon)$ by integrating, numerically or analytically, an appropriate renormalized integrand; rather than computing separately the quantities $I(\epsilon)$ and $I^{\rm ct}(\epsilon)$ and explicitly exposing the cancellation of the poles in $\epsilon$. For this purpose, a result such as (1) is not enough, as we will illustrate with simple examples below. On a more conceptual level, one may hope that understanding subtraction schemes at the level of the integrand may unearth geometrical structures underpinning the physics of renormalization, in analogy to how positive geometries have given a new perspective on the origin of locality and unitarity of scattering amplitudes - and simplified their computation Arkani-Hamed:2013jha ; Arkani-Hamed:2017tmz ; Arkani_Hamed_2018 ; arkanihamed2023loop .
@@
Here we propose a different strategy, based on the development of a suitable local subtraction scheme. We will explain how to build a collection of counterterms that reproduce the pole structure of an Euler integral, so that subtracting them from the original integrand yields a renormalized integrand that can be safely expanded under the sign of integration. Furthermore, the counterterm integrands can be integrated exactly, explicitly manifesting the pole structure of the Euler integral under consideration. In other words we look for a counterterm integrand $\mathcal{I}^{\rm ct}$ such that,
@@
$I(\epsilon)^{\rm ren}=\int_{0}^{1}\frac{dx}{x}\left[x^{\epsilon}(1+x)^{\epsilon}-x^{\epsilon}\right]=\int_{0}^{1}\frac{dx}{x}(\log(1+x)\epsilon+\dots)=\frac{1}{\epsilon}\frac{\zeta(2)}{2}+\dots,$ (11)
@@
Our first major result is a necessary and sufficient condition for a combination of Euler integrands to be locally finite, Th. 3. Building on this, we develop a simple-minded subtraction scheme, inspired by ideas discussed in Brown:2019wna in the context of string amplitudes, and already considered in the context of UV divergent Feynman integrals Hillman:2023ezp . These earlier works focus on specific instances of divergent Euler integrals: tree-level string amplitudes and UV divergent Feynman integrals, respectively. On the other hand, our subtraction scheme is applicable to any Euler integral that satisfies a certain geometrical property.
@@ 2 Polyhedral Geometry
A cone $\sigma$ in $\mathbb{R}^{n}$ is any set closed under taking positive spans. That is, if vectors $\{\rho_{i}\}_{i\in I}$ are in $\sigma$, then so is $\rm{Span}_{+}(\{\rho_{i}\}_{i\in I})$. A cone can be written as the positive span of a minimal collection of vectors called rays, which themselves cannot be written as the positive span of other elements of the cone. In other words, the rays $\mathrm{Rays\ }\sigma$ of a cone $\sigma$ form the minimal set such that
$\sigma=\mathrm{Span}_{+}\ \mathrm{Rays\ }\sigma.$ (17)
@@
$\mathrm{Poly}\left(\left(\begin{array}[]{c}0\\ 0\\ 4\\ 4\\ 0\\ \end{array}\right),\left(\begin{array}[]{ccc}-2&0&1\\ 0&-2&1\\ 2&0&1\\ 0&2&1\\ 0&0&1\\ \end{array}\right);(0),(0)\right),$
@@
$\mathcal{P}=\{1-\rho_{+}\cdot{\bf z}\geq 0\}\cap\{1-\rho_{-}\cdot{\bf z}\geq 0\}\cap\{1-\rho_{e}\cdot{\bf z}=0\},$ (27)
@@ 3 Newton Polytopes
Together, the value of $\mathrm{Trop\ }P(\rho)$ and the initial form $P|_{\rho}$ control the leading behavior of $P$ under the rescaling $\alpha\to\alpha\lambda^{-\rho}$, in the limit $\lambda\to 0$, which motivates us to understand them better. It turns out that they are respectively piecewise linear and piecewise constant on the fan $\Sigma=\rm{Fan\ Newt\ }P$. To see this, consider a cone $\sigma\in\Sigma$. For any $\rho$ in the interior $\sigma^{+}$ of the cone, the functional $\rho\cdot{\bf z}$ is always maximized over $\mathrm{Newt\ }P$ at the same face $F_{\sigma}$ of the Newton polytope. This shows that $P|_{\rho}$ is constant as $\rho$ varies in $\sigma^{+}$. Furthermore, over $\sigma$ we can compute $\mathrm{Trop\ }P(\rho)$ as $\rho\cdot{\bf m}$ for any ${\bf m}$ on the face $F_{\sigma}$,
@@
Consider the polynomial $P=1+x+y+xy$. Its Newton polytope and normal fan are shown in Fig.2 (a) and (b). We have $P|_{(-1,0)}|_{(0,-1)}=(1+y)|_{(0,-1)}=1=P|_{(0,-1)}|_{(-1,0)}$. On the other hand, $P|_{(-1,0)}|_{(1,0)}=1+y\neq x(1+y)=P|_{(1,0)}|_{(-1,0)}$.
@@ 4 Euler Integrals
At this point it should be clear that understanding the facet presentation of the Newton polytope of the Symanzik polynomials $\mathcal{U}$ and $\mathcal{F}$, that appear in the parametric form of Feynman integrals, is a mathematical problem of great physical interest. Currently, this presentation is not available if not in special cases Schultka:2018nrs . In particular, this means that there is no known purely graph theoretical algorithm - i.e. one that avoids solving a linear programming problem - to establish the convergence of an arbitrary Feynman integral.
@@ 4.2 Sector Decomposition
[…]
@@ 5 Tropicalization
$\rm{Trop\ }f(t)=\lim_{\alpha^{\prime}\to 0^{+}}\alpha^{\prime}\log|f(e^{-t/\alpha^{\prime}})|.$ (65)
If we identify $t$ with the logarithm of $\alpha$, we can think of the tropicalization of a function $f(\alpha)$ as capturing the asymptotic behavior of $f$ at the boundaries of the hyperoctant $\mathbb{R}_{+}^{n}$.
@@
where $\tau_{\lambda}$ is the operator returning the standard Taylor series expansion in $\lambda$. In practice, to compute $\tau_{\rho}$ one strips from the rescaled integrand $\mathcal{I}(\alpha\lambda^{-\rho})$ the overall power $\lambda^{-\mathrm{Trop\ }\mathcal{I}(\rho)}$, then performs a Taylor expansion on the rest and finally multiplies back in the factor $\lambda^{-\mathrm{Trop\ }\mathcal{I}(\rho)}$ (in complete analogy with the operators $\tau_{z}$ described in the Introduction). The structure of the Puiseaux series is illuminated by the notion of initial form of the integrand:
@@
Let us now consider again an Euler integrand $\mathcal{I}$ and two vectors $\rho_{1}$ and $\rho_{2}$ that are compatible according to the structure of the fan $\Sigma$ of the integrand, i.e. they belong to a common cone. Then we have
@@ 5.1 Local Finiteness for Euler Integrals
Note that while the proofs of the above theorems are based on the idea of decomposing the integrals into sectors, of which there is essentially one per vertex of $\mathrm{Newt\ }\mathcal{I}$, the result is that one can check a condition on each facet of $\mathrm{Newt\ }\mathcal{I}$. In general, the number of vertices and of facets of a polytope can be vastly different. For instance, in the case of Feynman integrals [^8] With some caveat regarding special kinematics already mentioned in 1 the relevant polytopes are generalized permutohedra Schultka:2018nrs , which have roughly $E!$ many vertices against $2^{E}$ many facets, $E$ being the number of edges of the graph. Therefore, it is practically useful to know about the above theorems, rather than performing sector decomposition and then apply Th. 1 to the resulting sector integrals.
@@ 6 Subtraction scheme
We anticipate that the geometric property is in general not satisfied by generic Euler integrals, nor by the specific instances that show up in physics applications. Therefore, this assumption currently represents the fundamental limitation of our subtraction scheme. However, note that any integral which does not satisfy the geometric property can always be expressed as a linear combination of integrals that do. This can be done algorithmically by performing a Nilsson-Passare analytical continuation along the rays for which the property does not hold.
@@
The combinatorics behind the above formula is familiar from the work of BPHZ on UV renormalization Hepp1966 ; Zimmermann1969 , and similar forest formulae have been written down also for IR divergences CHETYRKIN1982340 ; CHETYRKIN1984419 ; Herzog:2017bjx . The novelty compared with these previous results is that the renormalization map defined here produces an integrand that is not only finite but also locally finite. Furthermore, we will see shortly that our counter-terms can be integrated explicitly as many times as the order of the pole that they cancel, therefore manifesting the pole structure of the integral.
The following is the first main result of this section,
@@
Let us remark that any Euler integrand that does not satisfy the assumptions of Th. 6, either because it is power-divergent or because it does not satisfy the geometric property, can always be expressed as a collection of Euler integrands that do satisfy the assumptions. This can be algorithmically done by applying the Nilsson-Passare analytical continuation along the rays that break the assumptions above. Therefore, the subtraction formula can be applied to any Euler integral.
@@ 7.4 Double box
which agrees with the numerical values obtained by FIESTA Smirnov:2021rhf .
@@ 7.5 Eye Graph
$\Gamma(3\epsilon)^{-1}I(s;\epsilon)=\frac{1}{\epsilon^{2}}+\frac{7}{\epsilon}+(31-6\zeta(2))+\mathcal{O}(\epsilon),$ (113)
which is in agreement with previous results Chetyrkin:1981qh ; Hillman:2023ezp .
@@ 7.7 An angular integral
We provide another example beyond the case of Feynman integrals, a class of phase-space integrals known as angular integrals. In smirnov2024expansion it was introduced a parametric formulation which allows to apply our master formula to this case as well. In the language of smirnov2024expansion , we consider the four denominators, one-mass integral, given by
@@
The integrand has a power-divergence along the ray $\rho=(0,0,0,1)$, which requires analytic continuation in that direction to reduce log-divergent integrands. This produces two new integrands that satisfy the geometric property, hence we can apply our subtraction formula to them. Although the resulting integrands are not linearly reducible, they can still be integrated with HyperInt by rationalizing the last integration. In particular, the final result contains an interesting square root,
$r_{\pm}=-\frac{v_{14}v_{23}+v_{13}v_{24}-v_{12}v_{34}\pm\sqrt{(v_{14}v_{23}+v_{13}v_{24}-2v_{12}v_{34})^{2}+2v_{13}v_{23}(v_{12}v_{44}-2v_{14}v_{24})}}{2v_{13}v_{23}}.$
The result is given by
$I({\bf v};\epsilon)=\mathcal{N}\left(\frac{1}{\epsilon^{2}}J^{(-2)}+\frac{1}{\epsilon}J^{(-1)}\right)+\mathcal{O}(\epsilon),$ (122)
where
$J^{(-2)}=-\frac{v_{12}v_{14}v_{24}+v_{13}v_{14}v_{34}+v_{23}v_{24}v_{34}}{8v_{12}v_{13}v_{14}v_{23}v_{24}v_{34}},$ (123)
the expression for $J^{(-1)}$ involves polylogarithmic functions of weight two, and is reported in the ancillary files [^12] A mistake was done in the first version of this manuscript, and the reported result was wrong. We would like to thank V.Smirnov and F.Wunder for noticing the mistake. The analytic expression for this angular integral was previously unknown.